Walton Manufacturing Company (CMC) was started when it acquired $94,000 by issuing common stock. During the first year of operations, the company incurred specifically identifiable product costs (materials, labor, and overhead) amounting to $64,400. CMC also incurred $78,200 of engineering design and planning costs. There was a debate regarding how the design and planning costs should be classified. Advocates of Option 1 believe that the costs should be classified as general, selling, and administrative costs. Advocates of Option 2 believe it is more appropriate to classify the design and planning costs as product costs. During the year, CMC made 4,600 units of product and sold 4,000 units at a price of $38.00 each. All transactions were cash transactions.

Required

  1. a-1. Prepare an income statement and balance sheet under option 1.

  2. a-2. Prepare an income statement and balance sheet under option 2.

  3. b. Identify the option that results in financial statements that are more likely to leave a favorable impression on investors and creditors.

  4. c. Assume that CMC provides an incentive bonus to the company president equal to 14 percent of net income. Compute the amount of the bonus under each of the two options. Identify the option that provides the president with the higher bonus.

  5. d. Assume a 35 percent income tax rate. Determine the amount of income tax expense under each of the two options. Identify the option that minimizes the amount of the company’s income tax expense.
Answer #1

Ans) a1 & a2

Income statement Option 1 Option 2
Units sold 4000 4000
Sales (4000*$38) 152000 152000
less;Cost of Goods Sold( Unit product cost * units sold) 56000 124,000
Gross Margin 96000 28000
less; Selling & adm Exp 78200
Net Operating Income/Loss 17800 28000

Calculation;

Unit Product cost= Total Production cost/no of produced unit

Option1= 64,400/4600= 14

Option2= 64,400+78200/4600=31

Balance Sheet Option 1 Option 2
Assets
Cash (total Liability - Inventory) 103400 103400
Inventory (unsold units * Per unit product cost) 8400 18600
Total Current Assets 111800 122000
Total Assets 111800 122000
Liabilities
Common Stock 94000 94000
Retained Earnings 17800 28000
Total Liability & Equity 111800 122000

b) As the Net Operation Income is high in case of Option B, therefore it will leave a favorable impression on investors and creditors.

c) Option 2 provides higher amount of bonus.

Option 1 Option 2
Net Operating Income 17800 28000
Bonus @ 14% 2492

3920

d) Option 1 minimizes the amount of the company’s income tax expense.

Option 1 Option 2
Net Operating Income 17800 28000
tax @ 35% 6230 9800